### v2

# An Empirical Study of Training ID-Agnostic Multi-modal Sequential Recommenders

Abstract.
Sequential Recommendation (SR) aims to predict future user-item interactions based on historical interactions. While many SR approaches concentrate on user IDs and item IDs, the human perception of the world through multi-modal signals, like text and images, has inspired researchers to delve into constructing SR from multi-modal information without using IDs.
However, the complexity of multi-modal learning manifests in diverse feature extractors, fusion methods, and pre-trained models. Consequently, designing a simple and universal Multi-Modal Sequential Recommendation (MMSR) framework remains a formidable challenge. We systematically summarize the existing multi-modal related SR methods and distill the essence into four core components: visual encoder, text encoder, multimodal fusion module, and sequential architecture. Along these dimensions, we dissect the model designs, and answer the following sub-questions: First, we explore how to construct MMSR from scratch, ensuring its performance either on par with or exceeds existing SR methods without complex techniques. Second, we examine if MMSR can benefit from existing multi-modal pre-training paradigms. Third, we assess MMSR’s capability in tackling common challenges like cold start and domain transferring. Our experiment results across four real-world recommendation scenarios demonstrate the great potential ID-agnostic multi-modal sequential recommendation. Our framework can be found at: https://github.com/MMSR23/MMSR.
Multi-modality, Sequential Recommendation, Transfer Learning
copyright: acmcopyrightjournalyear: 2023doi: XXXXXXX.XXXXXXXprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
Figure 1. Sequential recommendation models the historical interactions between users and items and predicts the next item. Up: IDSR. Every item is represented by an ID Embedding. Bottom: MMSR. Every item is represented by its multi-modality information, such as covers and titles.
1. Introduction
Sequential Recommendation (SR) is a technique that leverages users’ past interactions to recommend their next item of interest. Its implementation in online environments, like e-commerce platforms and streaming media websites, has significantly expanded the range of options available to customers and has attracted increasing attention in the research community. One of the primary challenges in developing SR models is to generate high-quality user and item representations that can enhance recommendation performances.
In mainstream SR scenarios, a user can be represented by all the items that this user has acted upon, with each item encoded as a unique identifier (ID for short), as is shown in the upper part of Figure 1. To provide recommendation, the prevailing paradigm in current community is to maintain a learnable embedding matrix for all item IDs, and item representation in sequences can map as a latent vector by looking up its corresponding ID, we summarize this paradigm as ID-based Sequential Recommendation (IDSR). For example, GRU4Rec (Hidasi et al., 2015) adopts Recurrent Neural Network (RNN) as the sequence encoder, NextItNet (Yuan et al., 2019) adopts Convolutional Neural Network (CNN) as the sequence encoder, SASRec (Kang and McAuley, 2018) adopts unidirectional Transformer (Vaswani et al., 2017) as the sequence encoder, BERT4Rec (Sun et al., 2019) adopts bidirectional Transformer (Devlin et al., 2018) as the sequence encoder. Although straightforward and easy to implement, IDSR has several significant limitations. First, ID representations learned from one domain cannot be transferred across platforms, since the item vocabularies from two non-overlapping domains cannot be shared (Kang et al., 2020; Yuan et al., 2023). Second, IDSR can only learn the representations from the item IDs, leading to recommend cold items from highly skewed data distribution intractable (Wei et al., 2021; Zhu et al., 2021).
Alternatively, we can represent each item not by its ID, but by the multi-modality information (e.g., covers and titles) of the item, which we refer to as Multi-Modality based Sequential Recommendation (MMSR). Recently, several studies begin to utilize intrinsic modality or representations to glean item characteristics and learn user behaviors. For example, ZeroRec (Ding et al., 2021), UniSRec (Hou et al., 2022b), , VQRec (Hou et al., 2022a), MoRec (Yuan et al., 2023) and (Li et al., 2023a; Fu et al., 2023) pioneered the substitution of text or image representations for ID representations, achieving transferable sequential recommendations. TransRec (Wang et al., 2022) and MSM4SR (Zhang et al., 2023b) introduced multi-modal information to yield universal item and user representations. Compared with IDSR, MMSR is naturally endowed with strong transferability, since MMSR can process different items from different domains and platforms via the multi-modal contents of the items without the need for shareable item IDs. Furthermore, MMSR can alleviate the cold-start issue, since text and vision contents of the items provide richer semantics that are otherwise unobserved in the explicitly ID-based interactions.
We recap the basic idea of MMSR methods, which is to extract the multi-modal item embeddings from the pre-trained multi-modal item encoders ( an item encoder usually consists of a visual encoder, a textual encoder, and a multi-modal fusion module) and then adopt a sequence encoder (SASRec (Kang and McAuley, 2018)) to process the multi-modal item embeddings for recommendation training. On top of this backbone, these works design additional techniques to achieve better recommendation performance, such as enhancing the quality of text representations via parametric whitening and mixture-of-experts enhanced adaptor (Hou et al., 2022a), and achieving embedding alignment via contrastive learning with hard negatives (Hou et al., 2022b). Though effective, these techniques are not generalized enough to fit different types of multi-modal item encoders and sequence encoders and usually require careful tuning and selecting of hyperparameters. Meanwhile, the technical breakthroughs in NLP, CV, and RecSys are always evolving, with the new state-of-the-art text encoders (Devlin et al., 2018; Liu et al., 2019; Clark et al., 2020), vision encoders (He et al., 2016; Huang et al., 2020; He et al., 2022), and sequence encoders (Hidasi et al., 2015; Yuan et al., 2019; Kang and McAuley, 2018; Sun et al., 2019) emerging frequently. Excessive pursuit of the additional techniques leads to poor compatibility with different types of multi-modal item encoders and sequence encoders, making it difficult to utilize the technical breakthroughs in the NLP, CV, and RecSys communities.
To address this issue, this paper builds a simple and universal MMSR framework, which can support various types of multi-modal item encoders and sequence encoders and shows performance better or on par with the state-of-the-art multi-modal sequential recommenders. The core idea is to design a plug-and-play MMSR backbone consisting of multi-modal item encoders, multi-modal fusion modules, and the sequence encoder, such that different types of components can be seamlessly integrated into or replaced from the backbone. In our empirical study of MMSR, we try to answer the following key questions: Q1: How can we intuitively build MMSR from scratch, ensuring its performance on par with or even exceeds contemporary SR methods, all without the crutch of complex tricks? Q2: How can we build an MMSR framework such that it can directly benefit from the advances in vision-and-language communities? Q3: Will MMSR be supportive in addressing critical issues in the RecSys community, e.g., cold-start problems and cross-domain transferring?
To answer Q1, we conduct experiments on four large-scale industrial datasets under three SR architectures (i.e., SASRec (Kang and McAuley, 2018), GRURec (Hidasi et al., 2015), and NextItNet (Yuan et al., 2019)), and extensively test the performance of MMSR by equipping it with three TE, three VE and ten fusion strategies (including 4 vanilla coarse-grained, 4 enhanced coarse-grained and 2 fine-grained approaches). To answer Q2, we explore whether MMSR benefits from advances in existing Vision-Language Pre-trained (VLP) models by experimenting with different VLP architectures (e.g., CLIP (Radford et al., 2021) & GroupViT (Xu et al., 2022), ViLT (Kim et al., 2021) & VisualBERT (Li et al., 2019), and FLAVA (Singh et al., 2022)) as well as various representations of vision in VLP (i.e., region-based (Faster, 2015), grid-based (He et al., 2016) and patch-based (Radford et al., 2021)). To answer Q3, we put IDSR and MMSR in the cold-item scenario to evaluate their item representation ability and conduct transferring experiments on two source domains and two target domains, with the source and target datasets being scaled. In addition to exploring the above issues, practical guidance on how to effectively train MMSR would also be provided, e.g., whether full-finetuning the item encoders is necessary, what the impact of information volume brings (i.e., title length of sentences and mask ratio of images), and hyperparameter of different modules (see Appendix F). Our contributions can be outlined as follows:
  • To the best of our knowledge, we are the first to systematically explore multi-modal SR. Our observations not only thoroughly verify the effectiveness and universality of the proposed MMSR framework, but also provide much insightful guidance on developing more ID-agnostic recommendation models.
  • We conduct comprehensive empirical studies for MMSR along multiple dimensions, including text encoders (TE), vision encoders (VE), fusion strategies, multi-modal encoders paradigms, and RS architectures, which extensively validated the reliability of our results and provide diverse insights for this domain.
  • We bridge the divide between sequential recommendation and multi-modal learning communities, demonstrating how SR can leverage advancements in multi-modal learning to enhance recommendation quality.
  • We evaluate the performance of MMSR using datasets from different domains and platforms. Extensive experiments demonstrate its broad applicability for cold start and transferrable SR.
2. Related Works
2.1. ID-based Sequential Recommendation
Currently, the dominant paradigm of recommender systems is still IDSR, which assigns unique item/user IDs with one-hot embedding. There are thousands of recommenders based solely on IDs as input (Linden et al., 2003; Koren et al., 2009; Rendle, 2010; He et al., 2017; Hidasi et al., 2015). In this paper, we mainly focus on the most advanced and robust sequential recommenders, including RNN-based GRU4Rec (Hidasi et al., 2015), CNN-based NextItNet (Yuan et al., 2019), unidirectional Transformer-based SASRec (Kang and McAuley, 2018) and BERT4Rec (Sun et al., 2019). Although straightforward to implement, IDSR shows limited transferability and inferior performances under the cold-start setting.
2.2. Multi-Modal Sequential Recommendation
Some research focuses on improving recommendations by merging multi-modal content in IDSR. FDSA (Zhang et al., 2019) blends text with ID-embeddings using self-attention, and CARCA (Rashed et al., 2022) adopts cross-attention for item features. Yet, their reliance on the ID model limits cross-domain and platform flexibility.
To address this, ID-agnostic multi-modal recommenders have been introduced. UniSRec (Hou et al., 2022b) uses whitening (Su et al., 2021) to bridge the gap between item text and embeddings. VQRec (Hou et al., 2022a) matches item text with discrete index vectors to derive representations. MSM4SR (Zhang et al., 2023b) pioneers multi-modal information with contrastive learning for transfer. TransRec (Wang et al., 2022) leverages hybrid feedback for broader application. MoRec (Yuan et al., 2023) shows that text or vision can compete with ID representations. (Fu et al., 2023) further optimized MoRec, enhancing efficiency with Adapters (Rebuffi et al., 2017), reducing computational costs, and proving comparable performance to end-to-end methods. PMMRec (Li et al., 2023b) suggested an innovative cross-modal contrastive learning objective that enhances next-item prediction.
Our MMSR framework integrates diverse multi-modal item and sequence encoders, aiming to create a flexible, universally compatible, and efficient system. It outperforms traditional ID-based systems in addressing recommender system challenges like cold starts and transfer learning. A work closely related to ours is MMRec (Zhou et al., 2023), which outlines a clear process for multi-modal recommendation systems. However, it primarily targets GNN-based and general recommender systems, whereas our focus is on multi-modal sequential recommendation systems.
3. Framework
In this section, we introduce several essential modules in the MMSR framework and illustrate them in Figure 2. For an item with textual and visual information in SR, we fine-tune the corresponding pre-trained encoders to learn its text and vision features, followed by a fusion module to combine the two unimodal features into a cross-modal representation. Note, the combination of both text and vision encoders and fusion modules can be replaced by the existing multi-modal Vison-and-Language Pre-trained (VLP) models. Finally, we feed the item representation into a sequential recommendation network to complete the downstream recommendation task.
Figure 2. Illustration of MMSR. We systematically investigate MMSR along multiple dimensions, including vision and text encoders, multi-modality fusion, multi-modal pre-training paradigms, and downstream sequential architectures.
3.1. Text Encoder (TE)
In this paper, we take BERT (Devlin et al., 2018), RoBERTa (Liu et al., 2019) and ELECTRA (Clark et al., 2020) as the TE. There are two ways of using text embedding, i.e., either providing it to multiple text-specific transformers before the fusion or directly feeding it to a multi-modal fusion module. We adopt the former as our default setting, the latter will be discussed while introducing VLPs in Section 6.1.
3.2. Vision Encoder (VE)
Existing visual representations can be roughly divided into three categories: (1) Region Feature. Many traditional VLP models dominantly utilize region features, which are also known as bottom-up features. They are obtained from an off-the-shelf object detector such as Faster R-CNN (Faster, 2015). (2) Grid Feature. Generated by some CNN-based methods, e.g., ResNet50 (He et al., 2016), it avoids very slow region selection operations. (3) Patch or Patch-ViT. It is a commonly used approach that was first introduced by ViT (Dosovitskiy et al., 2020), showing outstanding performance in image classification tasks. In this paper, we first explore ResNet50 (He et al., 2016), MAE (He et al., 2022), and Swin (Liu et al., 2021) as VE in MMSR without VLPs, and then we give a detailed investigation of the three representations in VLPs in Section 6.2.
3.3. Multi-modal Fusion
As a key module that brings considerable influence in multi-modal architectures, fusion is also extensively studied in our MMSR. The fusion approaches we introduce can be divided into three classes: (1): Vanilla coarse-grained fusion methods including gate selection (Kiela et al., 2018), concatenation, summation, and FiLM (Perez et al., 2018). (2): Enhanced coarse-grained fusion methods via adding three-layer DNN to the vanilla ones, we make these modifications given that the original fusion methods may be too shallow for large text and vision encoders. (3): Fine-grained attention-based fusion approaches, including merged attention (Singh et al., 2022; Dou et al., 2022) and co-attention (Hendricks et al., 2021; Tan and Bansal, 2019). For the merged attention, linguistic and visual features are simply connected and fed to $1$ (or $n$) single-stream transformer block(s). In the co-attention module, two features are fed independently into different transformer blocks, followed by a merging block to force cross-modal interaction.
3.4. Vision-and-language Models (VLPs)
In this paper, SR is treated as a new downstream task of multi-modal communities. Here, we wonder if the existing VLPs paradigm can improve the task, the feasibility of which provides a theoretical potential for training specific large-scale multi-modal PTM for recommender systems.
The multi-modal community has developed some mature VLP models to get cross-modal semantics of text and vision information. We transfer these VLP paradigms to our MMSR to see their effect on recommendation tasks. Inspired by (Kim et al., 2021), VLPs can be split based on two points: (1) whether the two modalities have specific encoders; (2) whether the two modalities interact with each other in a deep fusion module. We first explore the impact of existing VLP architectures by directly introducing them into our framework, e.g., CLIP (Radford et al., 2021) & GroupViT (Xu et al., 2022), ViLT (Kim et al., 2021) & VisualBERT (Li et al., 2019) and FLAVA (Singh et al., 2022) (They stand for three typical paradigms of VLPs, i.e., two-stream and shallow fusion, single stream and deep fusion, two-stream and deep fusion), as is shown in Appendix E.
In our experiments, we note that variations in the architecture of vision models significantly affect the overall performances. Therefore, we carry out further exploration by evolving vision representation in VLPs. We select VisualBERT (Li et al., 2019) and replace the vision input with region-based (Faster-RCNN (Faster, 2015)), grid-based (ResNet50 (He et al., 2016)), and patch-based (CLIP-ViT (Radford et al., 2021)) visual representations in Section 6.2.
3.5. Recommendation architectures
We empirically select three emblematic SR architectures to evaluate MMSR, namely Transformer-based SASRec (Kang and McAuley, 2018), GRU-based GRURec (Li et al., 2022), and CNN-based NextItNet (Yuan et al., 2019). Specifically, we juxtapose SASRec, GRURec, and NextItNet to scrutinize the interplay between the most sequential recommendation architectures and the multi-modal item encoder. In order to ensure the efficiency of the model, we default use dense next-item prediction methodologies for all SR architectures inspired by (Kang and McAuley, 2018) (see Appendix B for details).
4. Experiments
4.1. Datasets
Table 1. Dataset statistics (after preprocessing).
Dataset #users #items #cold avg.length types
HM 200K 85,019 204,988 31.39 purchases
Amazon 100K 63,456 42,674 7.42 purchases
Bili 100K 76,178 5,632 17.95 comments
Kwai 200K 39,410 83,689 13.45 comments
Many existing modality related SR (Hou et al., 2022b, a) primarily focus on the e-commerce domain of item images with clean and pure backgrounds (e.g., items in the Amazon dataset all belong to the ”fashion” category). However, There is insufficient evidence on whether complex multimodal information can facilitate SR with complex content (e.g. short videos with broad topics). To provide a more comprehensive evaluation, we evaluate MMSR on four real-world datasets, i.e., the HM [^1] https://www.kaggle.com/competitions/h-and-m-personalized-fashion-recommendations/overview and Amazon [^2] https://nijianmo.github.io/amazon/index.html datasets for clothing purchase, the Bili [^3] https://www.bilibili.com/ and Kwai [^4] https://www.Kwai.com/?isHome=1 dataset (Zhang et al., 2023a; Cheng et al., 2023; Li et al., 2023b) from short-video platforms. For HM, we use descriptions and covers of the products and add categorical tags to the textual features to distinguish products with the same descriptions. For Bili and Kwai datasets, the title and cover of each video are regarded as the input. We consider those with less than 10 occurrences in the training set to select cold items in the cold scenarios. More details can be seen in Table 1 and Appendix A.
4.2. Default Implementation Details
[…]
4.3. Baselines
To provide a comprehensive evaluation on the performance of our method, we compare MMSR with the following baseline methods from three different groups [^5] We note that previous works seldom consider the multi-modal setting involving both texts and images. For a fair comparison, we select the strongest baselines (CARCA from group 2 and MoRec from group 3), and improve them into their multi-modal versions (denoted as CARCA++ and MoRec++).:
Pure ID-based sequential recommenders (IDSR) assign a unique ID to each item in the dataset.
  • GRURec (Hidasi et al., 2015) employs RNN as the sequence encoder and introduces ranking loss functions for model optimization.
  • NextItNet (Yuan et al., 2019) is a CNN-based sequential recommender, which combines masked filters with dilated convolutions to increase the receptive fields.
  • SASRec (Kang and McAuley, 2018) utilizes a unidirectional Transformer as the sequence encoder, which flexibly assigns attention weights to different items in the sequence.
ID-based sequential recommenders with side features (IDSR w. side feat.) enhance pure ID-based sequential recommenders with multi-modal side features such as texts and images.
  • FDSA (Zhang et al., 2019) captures the transition patterns between items and the contextual features of items (e.g., brands, text descriptions) via feature-based self-attention blocks.
  • CARCA++ is our improved version of CARCA (Rashed et al., 2022). The original version only supports imaged-based features, and we improve it into a multi-modal version supporting both texts and images.
Modality-based sequential recommenders learn transferable item representations via item texts or images.
  • VQRec (Hou et al., 2022a) maps item texts into discrete codes and then performs the lookup operation in the code embedding table to learn transferable item representations.
  • MoRec++ is our improved version of MoRec (Yuan et al., 2023). For a fair comparison, we improve it into a multi-modal version by fusing the representations extracted from text and the vision encoders.
4.4. Loss and Negative Sampling
For consistency in our evaluation, we employ the same loss function and sampling technique for MMSR and above baseline methods. Specifically, a negative sample is randomly selected (Yuan et al., 2023) and binary cross-entropy function as the loss. For each observed interaction $<u,j>$, it can be formulated as follows:
(6) $min-\sum_{<u,i,j>\in\mathcal{S}^{sample}}\left\{\log(\sigma(\hat{y}_{ui}))+\log(1-\sigma(\hat{y}_{uj}))\right\},$
where $\sigma(x)=1/(1+e^{-x})$.
4.5. Evaluation Metrics
We adopt the standard leave-one-out strategy(Kang and McAuley, 2018) (see Appendix C for explanations) and split the three used datasets into three parts: training, validation, and testing data, and evaluate all models in terms of two Top-N ranking metrics: Hit Ratio (HR@10) and Normalized Discounted Cumulative Gain (NDCG@10, or NG@10).
5. Performance of MMSR (Q1)
To systematically explore how to build MMSR from scratch to achieve the best results without using complex tricks, we conduct extensive experiments using (1) different text and vision encoders, (2) different multimodal fusion methods, and (3) different SR architectures. The best MMSR architecture is then comprehensively compared with current baselines.
5.1. Explorations of Various Item Encoders
To explore the potential of MMSR step by step, we first try different combinations of text and vision encoders without applying existing VLP models. Concretely, we explore the potential of our MMSR in unimodal mode [^6] Note that MMSR-T and MMSR-V in the tables are the ablation versions of our MMSR with only the text and vision information., then we construct the complete multi-modal MMSR by merging both TE and VE, i.e., we set RoBERTa (Liu et al., 2019) as the default TE and combine it with various VE, and choose MAE (He et al., 2022) as the default VE to integrate with other TE in reverse. For all of them, we select a one-block merge attention module as the fusion approach. All explorations are evaluated over HM and Bili datasets with SASRec, and the results are reported in Table 2.
Table 2. Comparisons of MMSR with different TE(i.e., BERT (Devlin et al., 2018), RoBERTa (Liu et al., 2019) and ELECTRA (Clark et al., 2020)) and VE (i.e., ResNet50 (He et al., 2016), MAE B-224/16 (He et al., 2022), and Swin B-224/16 (Liu et al., 2021)). All TE are in the base size. RoBERTa and MAE are default TE and VE, respectively.
- Encoders HM Bili
HR@10 NG@10 HR@10 NG@10
MMSR-T RoBERTa 8.2275 4.6923 3.7790 1.9655
ELECTRA 8.2010 4.6195 3.3690 1.7178
BERT 8.1810 4.6353 3.7990 1.9810
MMSR-V ResNet 7.3520 3.9049 3.2010 1.6550
MAE 6.1945 3.2329 2.4200 1.1823
Swin 7.8615 4.2045 3.3940 1.7281
MMSR Ro.+ResNet 8.4280 4.6649 3.6600 1.8942
Ro.+Swin 10.1095 5.7703 4.0810 2.2322
Ro.+MAE 9.9635 5.8046 4.2090 2.1194
ELE.+MAE 9.1675 5.1984 3.7940 1.9392
BERT+MAE 9.1100 5.1590 4.1300 2.1484
Various Text Encoders. There is no significant difference in performance for different TE in both MMSR-T and MMSR, while RoBERTa(Liu et al., 2019) always seems to achieve the best performance. This suggests that the capabilities of various TEs in feature representation have become very close, indicating that text representation techniques have matured considerably.
Various Vision Encoders. MAE (He et al., 2022) or Swin (Liu et al., 2021) perform the best in MMSR. However, the results of MAE (He et al., 2022) and ResNet (He et al., 2016)  [^7] Like PixelBERT (Huang et al., 2020), ResNet is used as the backbone to reduce the dimension of pixel features and we selected 49 vectors from each image. are exactly opposite. For the poor performance of ResNet (He et al., 2016) in MMSR, we guess it is because the current fusion approach cannot digest the outputs of ResNet (He et al., 2016) (CNN-based) and RoBERTa (Liu et al., 2019) (transformer-based) well. For MAE (He et al., 2022), the default high mask ratio (75%) may be the reason for its poor performance in MMSR-V.
Results. The performance of MMSR-T is generally better than that of MMSR-V, which may be that the dense semantics of textual information is easier to learn in SR (Hou et al., 2022b). The fully transformer-based (e.g., RoBERTa and MAE or Swin) perform the best in MMSR, which suggests the superiority of unified architectures.
Table 3. Comparison MMSR with various fusion methods. Coarse: MMSR with coarse fusion strategies. Coarse+: add a three-layer DNN based on the coarse fusion. CoAtten-1: co-attention fusion of 1 layer. MergeAttn-$n$: merge-attention fusion of $n$ layer.
- Fusion HM Bili
HR@10 NG@10 HR@10 NG@10
Coarse Gate 8.9505 5.0762 3.9120 2.0108
FiLM 8.6250 4.8787 3.8300 1.9809
Summation 6.9605 3.7458 3.4780 1.7709
Concatenation 6.7575 3.6387 3.4060 1.7318
Coarse+ Gate 8.4120 4.7515 3.5300 1.7819
FiLM 8.4760 4.7998 3.5340 1.7711
Summation 8.5270 4.8214 3.5960 1.8021
Concatenation 8.5035 4.8302 3.3690 1.6861
Attention CoAttn-1 7.6810 4.1932 3.8210 1.9463
MergeAttn-1 9.9635 5.7703 4.0810 2.1194
MergeAttn-2 10.0780 5.7448 3.8440 1.9844
MergeAttn-4 9.8975 5.7634 3.4210 1.6852
Table 4. Comparison of IDSR and MMSR with various SR architectures. HR@10 is as the default metric here.
Dataset Architecture IDSR MMSR
MMSR-T MMSR-V MMSR
Bili SASRec 3.3942 3.7790 3.3640 4.2090
GRURec 2.3247 2.9260 2.6400 3.2960
NextItNet 2.0521 2.3860 1.8920 2.7120
Kwai SASRec 4.7521 5.3110 4.7485 5.5260
GRURec 3.9487 4.2845 4.1315 4.7190
NextItNet 3.1538 3.7525 3.4195 4.6025
HM SASRec 8.4672 8.2670 7.8615 10.1095
GRURec 6.1241 5.2780 5.7215 6.7965
NextItNet 6.1752 6.5275 5.8635 7.7755
Amazon SASRec 19.1250 19.2436 19.1529 20.1095
GRURec 16.0417 16.0056 15.0926 18.3956
NextItNet 15.2378 16.2342 15.9076 17.7984
Table 5. Performance comparison (%) on four datasets. The best and the second-best performance in each row are bolded and underlined, respectively. Improvements compared with the best baseline method are indicated in the last column.
Dataset Metrics IDSR IDSR w. Side Feat. Modality-based SR Ours Improv.
GRURec NextItNet SASRec FDSA CARCA++ VQRec MoRec++ MMSR
Bili HR@10 2.3247 2.0521 3.3942 3.7623 4.0456 1.3569 3.7531 4.2210 +4.16%
HR@20 3.3678 3.1256 4.4350 4.9143 6.1476 2.0721 5.8824 6.3135 +2.63%
HR@50 6.0234 5.1288 7.9087 8.9260 9.9689 3.5731 8.8856 10.4437 +5.29%
NDCG@10 1.2145 1.0378 1.6729 1.7934 2.1159 0.6089 1.9834 2.2315 +5.17%
NDCG@20 1.4271 1.2876 2.0174 2.2837 2.5765 0.7167 2.4829 2.7046 +4.74%
NDCG@50 1.9463 1.8291 2.6175 2.9750 3.3290 1.0827 3.2254 3.5287 +5.66%
Kwai HR@10 3.9487 3.1538 4.7521 4.9357 4.9571 1.9523 4.9518 5.3786 +7.84%
HR@20 5.9744 5.1966 6.8462 6.9929 7.1286 2.9568 6.7517 7.7794 +8.37%
HR@50 9.8718 8.6068 10.9051 10.9714 11.3502 4.8528 11.7659 12.3295 +4.57%
NDCG@10 2.0598 1.9915 2.5043 2.6664 2.5857 0.8732 2.6297 2.8576 +7.99%
NDCG@20 2.5641 2.2137 3.0513 2.8317 3.0857 1.2083 3.1294 3.4085 +8.19%
NDCG@50 3.3419 2.9744 3.8556 3.8956 4.0652 1.5093 4.0086 4.3092 +5.66%
HM HR@10 6.1241 6.1752 8.4672 8.8214 9.7678 4.1623 9.6933 10.0425 +2.82%
HR@20 8.3577 8.3066 11.0511 11.2213 12.1035 5.7416 11.9867 12.6694 +4.47%
HR@50 12.6058 12.7372 15.5547 15.4623 16.6235 8.0028 16.0933 17.5643 +5.36%
NDCG@10 3.6350 3.5328 5.4715 5.6917 6.4234 2.2235 6.1400 6.3646 -0.91%
NDCG@20 4.1898 4.0431 6.1153 6.4784 6.8217 2.6315 7.0533 6.9224 +1.46%
NDCG@50 5.0292 5.0146 7.0146 6.6347 7.5158 3.0562 7.0867 7.8824 +4.86%
Amazon HR@10 16.0417 15.2378 19.1250 18.7667 19.7250 17.7143 19.2512 19.6412 -0.42%
HR@20 17.6256 15.9234 19.3078 18.8245 19.8567 19.0567 19.5923 19.9934 +0.68%
HR@50 19.3023 17.3445 20.9645 19.4423 21.3012 20.4012 21.2578 21.9023 +2.81%
NDCG@10 14.9923 12.9945 16.7067 15.8512 17.1434 12.8034 17.0767 17.3656 +1.30%
NDCG@20 15.2412 13.2323 16.9345 16.3145 17.8312 13.0434 17.9245 18.0201 +1.20%
NDCG@50 15.5745 13.5012 17.2312 17.0278 18.5145 13.3923 18.3934 18.6512 +0.77%
5.2. Impact of Various Fusion Approaches
In this section, we show investigations on various multi-modal fusion strategies for MMSR with a default of TE, VE and RS architecture (i.e., RoBERTa (Liu et al., 2019), MAE (He et al., 2022), and SASRec (Kang and McAuley, 2018)). Specifically, we explore four vanilla coarse-grained fusion methods (i.e., gate selection (Kiela et al., 2018), concatenation, summation, and FiLM (Perez et al., 2018)), four enhanced coarse-grained fusion methods (with a three-layer DNN applied to the previous ones), and two fine-grained fusion methods (i.e., co-attention (Hendricks et al., 2021) and merge-attention (Dou et al., 2022)).
Coarse-grained fusion. As shown in Table 3, MMSR with coarse-grained fusion lacks competitiveness, especially in the HM dataset. Among them, gate (Kiela et al., 2018) and FiLM (Perez et al., 2018) do slight favor to MMSR while summation and concatenation seriously deteriorate the performance. Comparing Coarse with Coarse++, we notice that added DNN drives all the shallow fusions to similar performance. It seems the three-layer DNN overwhelms the multi-modal fusion.
Fine-grained fusion. The results in Table 3 show that merge-attention outperforms co-attention. We also find that there is no improvement even though we deepen the merge-attention from one to four transformer blocks, and the results on the Bili dataset are even inversely proportional to the depth. It may be that the amount of data is not enough to support the learning of deep fusion. It suggests that VLP models should be considered if we want to use deeper fusion modules (see Section 6.1).
Results. Coarse-grained fusion does not guarantee the performance of MMSR. As for fine-grained methods, merge-attention exceeds co-attention in integrating different modalities.
5.3. MMSR with Various SR Architectures
Based on the above explorations, we choose powerful RoBERTa (Liu et al., 2019) and Swin (Liu et al., 2021) as text and vision encoders, respectively. To validate the robustness of our MMSR, we consider the effect of various architectures, including SASRec (Kang and McAuley, 2018), GRURec (Hidasi et al., 2015), and NextItNet (Wang et al., 2022), as shown in Figure 2 (right). We evaluate the performance over four recommendation datasets (i.e., HM, Bili, Kwai, and Amazon).
Results. As shown in Table 4, MMSR configured with SASRec shows strong competitiveness. Furthermore, no matter which SR architecture is given, the performance of traditional ID-based SR (IDSR) are much lower than that of MMSR. The ablation study shows that single-modal information, is still very competitive, even surpassing IDSR in some cases. This suggests that current CV and NLP techniques have enabled id-agnostic recommendation models to excel, highlighting the potential of the MMSR paradigm.
5.4. Overall Performance Comparisons
We evaluate MMSR and other baseline methods on 4 source datasets. From the results of Table 5, we can reach the following conclusions:
  • Against other baselines, MMSR consistently comes out on top in most tests. What’s more, it performs just as well, if not better than, methods that rely on IDs, even those with added side features. This suggests that MMSR can make sense of multi-modal data for recommendations without needing item IDs.
  • Compared with the best baseline CARCA++, MMSR shows smaller performance gains on HM and Amazon datasets, but larger gains on Bili and Kwai datasets. This is probably because the former two datasets have relatively clean and pure image backgrounds, while the latter has more complex visual presentations (e.g., posters). Benefiting from the powerful and robust backbone derived from the above sufficient exploration strategy, MMSR can effectively cope with data noise and improve performance, showing greater performance improvements on the Bili and Kwai datasets.
  • Our experiments show that MoRec and our proposed method outperform VQRec [^8] We observe VQRec do not perform well under our experimental setting, which might be attributed to the following reasons: non-end-to-end training manner (Yuan et al., 2023), lack of multimodal information., indicating that end-to-end learning is more effective than traditional pre-extraction. This is likely due to the fact that end-to-end learning allows the item encoder to better capture the semantic information of modalities (Yuan et al., 2023).
6. Inherit Advances in Multi-modal Pre-training Community (Q2)
MMSR employs SR as a subsequent task in multimodal learning, prompting us to examine if advancements in the multi-modal pre-training community, i.e., Vision-and-Language Pre-training (VLP), can enhance MMSR. In particular, we delve into two fundamental questions frequently addressed in the VLP community: (1) the architecture of VLP model (Singh et al., 2022; Kim et al., 2021), and (2) the visual representation within VLP model (Kim et al., 2021).
Table 6. Comparison of MMSR with various VLP models of different architectures. (a): two-stream & shallow fusion. (b): single stream & deep fusion. (c): two-stream & deep fusion.
VLPs HM Kwai
HR@10 NG@10 HR@10 NG@10
a CLIP 6.4240 3.3109 4.2510 2.1332
GroupViT 7.7935 4.2868 4.1800 2.1132
b ViLT 9.3465 5.3226 4.9540 2.5938
VisualBERT 9.2680 5.2937 5.0740 2.6141
c FLAVA 10.1435 6.5820 5.4040 2.7125
Table 7. Comparison of vision representations for MMSR.
- Representations HM Kwai
HR@10 NG@10 HR@10 NG@10
IDSR - 8.4672 5.4715 4.7521 2.5043
VisualBERT Region 9.2680 5.2937 5.0740 2.6141
Grid 9.5875 5.5167 5.2250 2.6814
Patch-Emb 9.0205 5.1664 5.0695 2.6255
Patch-ViT 9.8525 5.7026 5.2502 2.6374
6.1. Impact of Various VLP Architectures
Our exploration initiates with a meticulous examination of diverse VLP architectures. To be specific, these can be broadly classified into three emblematic paradigms: (1) Two-stream and shallow fusion, as exemplified by models such as CLIP (Radford et al., 2021) and GroupViT (Xu et al., 2022), (2) Single-stream and deep fusion, represented by the likes of ViLT (Kim et al., 2021) and VisualBERT (Li et al., 2019), and (3) Two-stream and deep fusion, encapsulated by models like FLAVA (Singh et al., 2022).
For architectures such as CLIP (Radford et al., 2021) and GroupViT (Xu et al., 2022) that lack a fusion module, we elected to employ the widely-adopted concatenation technique to glean cross-modal representations. For all transformer-based encoders, we limited our fine-tuning to the two top-layer transform blocks, thereby adhering to our prior experiments. In the case of FLAVA, considering its substantially deep fusion module, we only fine-tune its final fusion block.
Results. In Table 6, we can see that MMSR performs worst in the CLIP & GroupViT architecture. By analyzing in conjunction with the other two VLPs, we hypothesize it is the shallow fusion that deteriorates the performance. We also notice that FLAVA drives MMSR to completely surpass other VLPs, which is consistent with its excellent performance in the multi-modal fields. Note, the best-performing MMSR in RQ(1) has a similar architecture to FLAVA, which may imply that the success of PTM in other domains can be transferred to the recommendation task.
6.2. Impact of Various Vision Representation
We carry out further exploration by evolving vision representation in VLPs (see Section 3 for details). We select VisualBERT (Li et al., 2019) as our foundational model due to its iconic status within the domain of VLPs. Most subsequent strides of vision representation are indeed enhancements built upon it. To validate the MMSR, we have incorporated an assortment of vision representations, including region-based (Faster-RCNN (Faster, 2015)), grid-based (ResNet50 (He et al., 2016)), patch-Embedding (CLIP-ViT (Radford et al., 2021), excluding 12 transformers) and patch-Vision Transformer (CLIP-ViT, including 12 transformers). These representations typify a significant trajectory in the evolution of VLPs, thereby prompting our curiosity to investigate their adaptability to MMSR.
Results. As is shown in Table 7, all MMSRs with various vision representations can exceed IDSR, and patch-ViT representation is the best, which is in line with the development process in the field of VLPs. Region-based and Patch-Emd representations underperform, whereas MMSR with Grid-based and Patch-ViT demonstrate notable improvements over IDSR. This disparity might be attributed to the strong augmentation of the latter two specific vision encoders, namely ResNet50 (He et al., 2016) and CLIP-ViT (Radford et al., 2021). Note that in Section 5.1, ResNet50 (grid-based) performs poorly in the previous setting. However, it achieves good results here, suggesting that deep fusion may resolve instability caused by inconsistent representations.
7. Cold-start and Transferability (Q3)
7.1. MMSR in Cold-start Scenario
In this section, we concentrate on the cold-start problem. MMSR is naturally suitable for such scenarios since their encoders are especially suitable for modeling the raw features of items, i.e., text and image, whether cold or not. To verify this, we conduct evaluations using IDSR, MMSR-T and MMSR-V, and MMSR in the cold-item scenario of the above 4 datasets.
Table 8. Performance comparison (%) under cold-start setting.
Dataset Metrics SASRec MMSR-T MMSR-V MMSR
Bili HR@10 0.1192 1.5493 0.9296 1.3824
NDCG@10 0.0450 0.7722 0.4706 0.7603
Kwai HR@10 0.0419 3.9812 3.9408 4.7393
NDCG@10 0.0450 2.0438 1.9730 2.3555
HM HR@10 0.0778 2.9385 1.8755 2.7523
NDCG@10 0.0406 1.7457 1.0095 1.5740
Amazon HR@10 0.1723 4.9190 4.4885 5.6222
NDCG@10 0.1381 4.8547 4.3779 5.4046
Results. Table 8 reveals that MMSR, MMSR-T, and MMSR-V all notably outperform IDSR in predicting cold items. This superior performance stems from the fact that IDSR only derives representations from behavioral patterns, but MMSR incorporates original modal features, the advantage of which becomes more apparent for cold start. Interestingly, while MMSR-T and MMSR-V have comparable performances in standard situations (as is shown in Table 4), MMSR-T tends to be more adept at handling cold items. This suggests that text-based representations might have a stronger generalizing capability for recommendation tasks than visual ones.
7.2. Results of Transfer Learning
We conduct the transfer learning experiments only on MMSR since IDSR is non-transferable.
Scaling the source domain. As shown in Table 9, we choose the HM-100K, -200K, and -500K datasets as the source domain and the e-commerce Amazon-10K (He and McAuley, 2016; McAuley et al., 2015) dataset as the target domain. As the scale of the source domain increases, the results of transferring increase compared with those from scratch. This implies that by increasing the sample size, MMSR can more effectively uncover user preferences within the modal content.
Scaling the target domain. On the contrary, we select Kwai-200K as the source dataset and scale the target Bili dataset from 10K, 20K to 50K. As shown in Table 10, the transferring improvement is reduced as the target domain expands.
Results. As is shown in Table 9 and Table 10, MMSR can perform better than training from scratch, but it does not always hold when the ratio of the source domain to the target domain is narrowed, e.g., the results of transferring from Kwai-200K to Bili-20K and -50K show no advantages. The performance drop might stem from differences in user groups between domains. Larger downstream datasets can introduce unrelated features and noise, hampering the transfer learning performances of sequential recommenders.
Table 9. Transfer learning from HM to Amazon. Source: HM-100K, HM-200K, HM-500K. Target: Amazon-10K Clothes Dataset. ‘Improv.’ is the relative improvement of the best transferring compared to training from scratch.
Metrics IDSR Scratch Transfer Improv.
100K 200K 500K
HR@10 20.2600 25.4900 26.3100 26.2400 26.4900 +3.80%
NDCG@10 17.2302 21.7594 23.3476 22.9698 23.3558 +6.83%
Table 10. Transfer learning from Kwai to Bili. Source: Kwai-200K dataset. Target: Bili-10K, -20K, and -50K dataset. ‘Improv.’ is the relative improvement of the transferring compared to the corresponding training from scratch.
10K 20K 50K
HR@10 NDCG@10 HR@10 NDCG@10 HR@10 NDCG@10
ID 1.0300 0.5252 1.3950 0.7633 2.0420 1.0727
Scratch 1.4600 0.7087 2.0750 1.1148 2.7580 1.3801
Transfer 1.5600 0.8101 2.1600 1.0587 2.7760 1.3689
Improv. +6.41% +14.31% +3.93% -5.03% +0.64% -0.81%
8. Conclusion
In this paper, we propose a simple and universal Multi-Modal Sequential Recommender (MMSR) framework. We conduct comprehensive empirical studies along multiple dimensions, including text encoders, vision encoders, fusion strategies, multi-modal pre-training paradigms, and SR architectures. The experimental results demonstrate that MMSR not only completely outperforms the traditional IDSR, but also own decent transferability across platforms. Moreover, our work suggests an empirical potential for training specific multi-modal PTMs for unified SR. As mentioned above, this paper is only a preliminary study of MMSR with the following limitations: (1) we only consider SR scenarios with text and visual modalities, not voice and video; (2) we only consider a very basic end-to-end way to introduce TE, VE, and VLPs into the recommendation model, without introducing any other technical tricks, the results we obtain from MMSR may still be suboptimal. To further serve RecSys communities, we plan to release a large-scale multi-modal RecSys dataset in the future, and specific multi-modal PTMs for unified SR frameworks.
Appendix A Details of Datasets
The HM and Amazon of clothes category dataset includes customer-product (user-item) interactions and customer and product metadata, e.g., category and color, descriptions, and covers, as shown in Figure 3 (left). In this paper, we only use texts and images, adding some category information as part of the description to further distinguish different items. Following (Yuan et al., 2019), we removed users with less than 5 interactions simply, and randomly selected 200K and 100K users.
To build the Bili dataset, we randomly crawl videos (less than 10 minutes in duration) of different categories from the online platform. We record the public comment information of these videos as interactions (i.e., we think the user would be interested in this video if a user comments on a video). We do not capture any user’s personal information about privacy, whose IDs are already anonymous after data masking. The video metadata such as cover and title are recorded as shown in Figure 3 (right). The Kwai dataset is collected similarly to the Bili dataset. Then we also removed users with less than 5 interactions and randomly selected about 100K and 200K users from the Bili and Kwai datasets, respectively.
We take some basic pre-processing by setting the size of all images to 224 × 224, with a maximum of 50 words for all text descriptions or titles (covering more than 95% of the descriptions). Noted, the datasets used in this paper should ensure that the user’s decision (i.e., whether to interact with the item or not is determined only by the modality characteristics of the item). But in fact, none of the 2 datasets meets this assumption. For example, except for the textual descriptions and cover of the products, users’ purchase decisions on the HM dataset are also influenced by the price of the item. However, we can still achieve better performance even taking only text and image representations of items, which proves the potential of MMSR over IDSR.
Besides, for all datasets, we consider cold scenarios (different from regular scenarios). That is, we count the interactions of all items in the training set and consider those items with less than 10 occurrences as cold items (Anonymous, 2023). Then, we truncate the complete user sequence to get some sub-sequences with a cold item at the end, all of which are used to evaluate the capability of MMSR in the cold scenario.
Appendix B Details of Various SR Architectures
In this study, we undertake a comprehensive empirical analysis using three representative SR architectures to rigorously assess the efficacy of MMSR. The selected architectures encompass a broad spectrum of methodologies, specifically: the Transformer-based SASRec (Kang and McAuley, 2018), the GRU-based GRURec (Li et al., 2022), and the CNN-based NextItNet (Yuan et al., 2019). Our main objective is to explore the dynamic interrelationship between these prevalent sequential recommendation architectures and the intricacies of the multi-modal item encoder. To maintain consistent efficiency, we adopt dense next-item prediction methodologies for all these SR architectures(Kang and McAuley, 2018).
(3) $R(x)=\prod\limits_{i=0}^{l}r(x_{i}|x_{0:l-1},\theta)r(x_{0})$
where $r(x_{i}|x_{0:l-1},\theta)$ means the probability of $i$-th item conditioned on all previous item $x_{0:i-1}$.
Obviously, joint distributions are more efficient at capturing sets of all sequence relations, and transformers are more powerful than GRU and CNN basic modules. Hence, GRURec and NextItNet would be taken as degenerated versions of SASRec. Our results on the above 3 SR architectures can prove that MMSR is widely applicable in most cases. We show some specific hyperparameters for SASRec, GRU4Rec, and NextItNet. The layer number is set to 2, 4, and 2, respectively, which is from searching in $\{1,2,4,8\}$. The attention headers of SASRec are set to 2.
Appendix C Evaluations Metrics
In this paper, by employing a standard ”leave-one-out” strategy (Kang and McAuley, 2018), we split the dataset into training, validation, and test sets. For example, we selected the latest 23 items of each user sequence for all datasets here (i.e., 1-21 items for training, 2-22 for validation, and 3-23 for testing).
We evaluate all models using two Top-N ranking metrics: HR@N (Hits Ratio) and NDCG@N (Normalized Discounted Cumulative Gain) (Yuan et al., 2019), where N is set to 10 in this paper. Specifically,
(4) $HR=\frac{1}{N}\sum_{i=0}^{N}hits(i)$
$hits(i)$ predict whether the user’s $i$-th intention item is in the recommended rank list (1 if it is, and 0 otherwise).
(5) $NDCG=\frac{1}{N}\sum_{i=0}^{N}\frac{1}{log_{2}(p_{i}+1)}$
$p_{i}$ means the position of the user’s $i$-th intended item in the recommendation list. It would be close to +$\infty$+ ∞ if there is no user’s $i$-th intended item in the recommendation rank list.
Following (Krichene and Rendle, 2022), we rank the ground-truth target item by comparing it with all the other items in the item pool. Finally, we report the results of the testing set by choosing checkpoints with the best validation performance.
Appendix D Performance of MMSR (Q1)
Figure 4 shows the convergence of IDSR, MMSR-T, MMSR-V, and MMSR with the best setting in Table 2 and 4. i.e., SASRec RS architecture, RoBERTa (Liu et al., 2019) and Swin B-224/16 (Liu et al., 2021) as default text and vision encoder, respectively.
Figure 4. Convergence of IDSR, MMSR-T, MMSR-V and MMSR on HM, Bili, and Kwai datasets.
It can be seen that there is a performance gap between the validation and the test results by comparing Figure 4 and Table 4. In many validation experiments, MMSR-T and MMSR-V are not comparable to IDSR, but they are comparable or even surpassed IDSR on the test set, e.g., the HM-200K and Kwai-200K datasets. This implies the better generalization performance of the modality-based recommendation than IDSR. Especially, MMSR has always been able to greatly surpass IDSR, no matter how on the validation set and on the test set.
Appendix E MMSR with Various VLPs (Q(2))
In this paper, we introduce three typical VLP paradigms to our MMSR to see their impact on the recommendation task, i.e., CLIP (Kang et al., 2020) and GroupViT (Xu et al., 2022) (two-stream and shallow fusion), VsualBERT (Li et al., 2019) and ViLT (Kim et al., 2021) (single stream and deep fusion), and FLAVA (Singh et al., 2022) (two-stream and deep fusion), as shown in 5. We show the convergence of MMSR with the above VLPs in Figure 6.
Figure 5. Description of each module: a) CLIP-B-224/32 and GroupViT (a CLIP variant for semantic segmentation); b) VisualBERT and ViLT; c) FLAVA.
Specifically, we choose SASRec as the default RS architecture. For CLIP and GroupViT, we fine-tune two transformer blocks for the text and vision encoders and use a concatenation layer to connect both as the input of the downstream recommendation network. Since VisualBERT can not extract vision features itself, we use Faster-RCNN(Faster, 2015) with a threshold of 0.05 and select 49 regions from each image as visual input. The last 4 transformer blocks of VisualBERT are fine-tuned, which is consistent with ViLT. As for FLAVA, we fine-tune the top two transformer blocks of text and vision encoders and all fusion modules.
Figure 6. Convergence of MMSR with various VLPs of 3 different architectures.
Figure 6 demonstrates MMSR with CLIP or GroupViT performs the worst. A possible hypothesis is that shallow fusion leads to performance degradation. We note that FLAVA drives MMSR to completely outperform IDSR, which is consistent with its excellent performance in the multimodal community.
E.1. Transferring of MMSR (Q(3))
Figure 7 show the convergence of the transferring experiments across platforms of MMSR. We take HM-500K and Kwai-200K as source domain datasets, and Amazon-10K and Bili-10K as target domain datasets, respectively (see Table 9 and 10). We train IDSR for 300 epochs, and the learning rate decays linearly to 10% of the original learning rate after 270 epochs. And both MMSR-Transfer and MMSR-Scratch are only trained for 60 epochs (already converged), and the learning rate decays linearly to 10% after 30 epochs. Compared with MMSR-Scratch, MMSR-Transfer achieves surprising results, i.e., more stable, faster, and better performance.
Figure 7. Comparison of MMSR-Transfer and MMSR-Scratch. Note IDSR is trained for 300 epochs, and we only record results every 5 epochs for it.
Appendix F Bag of Tricks
In this section, we provide some practical insights and helpful tricks for training a performant MMSR. If not specified, we select a default setting of TE, VE, SR architecture, and dataset (i.e., RoBERTa, MAE, SASRec, and Bili).
F.1. Full Fine-tuning vs. Top-layer Fine-tuning
In all the above experiments, we fine-tune only the top blocks for all encoders by default, which is mainly due to computational consumption. Here, we try to investigate whether all parameters need to be adapted. Table 11 shows the results of full and top-layer fine-tuning showed irregular fluctuations. Among them, top-layer fine-tuning works better in MMSR. We conclude that top-layer fine-tuning is worth considering in order to balance accuracy and efficiency in practice.
Table 11. Full vs top-layer fine-tuning.
- Encoders HM Bili
HR@10 NG@10 HR@10 NG@10
ID - 8.4825 5.2432 3.1670 1.7031
MMSR-T RoBERTa 8.2275 4.6923 3.7790 1.9655
RoBERTa-Full 8.6835 5.0220 3.7410 1.9294
MMSR-V MAE 6.1945 3.2329 2.4200 1.1823
MAE-Full 7.3975 4.0094 3.1410 1.6026
ResNet 7.3520 3.9049 3.2010 1.6545
ResNet-Full 7.3810 3.9253 3.1060 1.6019
MMSR Ro.+MAE 9.9635 5.7703 4.0810 2.1194
Ro.-F+MAE-F 9.5235 5.4953 4.0500 2.1063
F.2. Different Learning Rates
As shown in Table 12, using the same learning rate for all modules of the model will lead to performance degradation, which may be because the pre-trained parameters already contain a certain amount of knowledge about vision and language, and over-tuning them may lead to loss of this valuable information. As shown in Table 12, the performance of MMSR would deteriorate when all modules use the same learning rate. This is probably because the pre-trained parameters in the item encoders already contain a certain amount of knowledge about vision and language, and over-tuning may lead to the loss of this knowledge. But the downstream recommendation networks and fusion modules need training from scratch, so a large learning rate would be better.
Table 12. Comparison of MMSR with different learning rates for different modules. Top: TE and VE of MMSR. Downstream: the recommendation networks and fusion modules.
Learning rate MMSR
Top Downstream HR@10 NDCG@10
5e-5 5e-5 3.8300 1.9996
8e-5 8e-5 4.0400 2.1171
1e-4 1e-4 4.0810 2.1194
8e-5 1e-4 4.0500 2.1096
5e-5 1e-4 4.0910 2.1331
F.3. Length of Texts
In our investigation, we dedicated substantial effort to analyzing the impact of varying quantities of textual information. This aspect is crucial, as the volume of textual data in a recommendation system can significantly influence its predictive accuracy. Our primary focus was on discerning how changes in the length of titles, an integral component of textual data, affected the performance of our models. As demonstrated in Table 13, our exploration yielded noteworthy results. Both MMSR-T and MMSR exhibited a positive response to increases in title length. This trend implies that an expansion of title content can contribute to enhanced model performance. Our findings underline the significance of textual volume, particularly in title length, as a determinant factor for improving the functionality of MMSR.
Table 13. The impact of various lengths of text for MMSR-T and MMSR.
Tokens MMSR-T MMSR
HR@10 NDCG@10 HR@10 NDCG@10
10 3.5360 1.8597 3.6710 1.9008
30 3.7650 1.9455 3.9810 2.0765
50 3.7790 1.9655 4.0810 2.1194
F.4. Mask Ratio of Images
The volume of textual and visual information can greatly affect model performance MMSR. Inspired by MAE (He et al., 2022), we control the image volume by adjusting its mask ratio. Table 14 shows that as the mask ratio increases, the performance of MMSR-V rises but MMSR decreases gradually. A possible reason is that information is insufficient for MMSR-V thus it would be better to keep more visual features. While for MMSR, the information may be saturated with both linguistic and visual features being fed, and it would be more intractable to optimize since the redundant image information would be taken as noise in this case.
Table 14. The impact of the volume of image information for MMSR. Note that we use mask ratio from MAE (He et al., 2022) to control the volume of image information.
Mask Ratio Vision MMSR
HR@10 NDCG@10 HR@10 NDCG@10
0% 2.9130 1.4704 3.7990 1.9756
50% 2.6500 1.2560 3.9780 2.0996
75% 2.4200 1.1823 4.0810 2.1194
F.5. Visualization of IDSR and MMSR
In this study, we approach SR as a unique downstream application of multi-modal learning. Specifically, we employ solely multi-modal representations to ascertain whether they could match or even supersede traditional item-ID representations. We visualize both MMSR and IDSR from two perspectives of item and user. For this purpose, we selected the Bili dataset, randomly choosing 1000 items and 10 users, and then leveraged t-SNE (Van der Maaten and Hinton, 2008) tool to unveil their latent distributions.
F.5.1. Visualization of Items
In our study, we examine the visual representation of items using item-ID, as presented in Figure 8. The insights gathered from this analysis revealed two prominent characteristics. Firstly, we observed a notable clustering of popular items. These items, having a high frequency of interaction, tend to dominate and overshadow the less prevalent items in the dataset. Secondly, we noticed that ’cold’ items, which are items with few interactions, tend to be scattered and emerge as outliers. These observed tendencies are problematic as they can impede the recommendation system’s learning efficacy. Specifically, the recommendation model may over-emphasize popular items, creating a bias, while simultaneously struggling to learn or incorporate ’cold’ items due to their sparse presence. This disparity can lead to a reduced overall quality of recommendations, especially for users with niche tastes.
On the other hand, our study also highlights the potential advantages of multi-modal representations. Notably, these representations exhibit a balanced distribution of items, showing little sensitivity to popularity variations. The key to this resilience lies in the multi-modal representation’s ability to tap into the semantic richness of an item’s text and imagery. By leveraging these additional data dimensions, multi-modal representations effectively mitigate the ’overheating’ phenomenon associated with popular items and counteract the ’cold’ issue linked with less interacted items.
Figure 8. Visualization of items in IDSR and MMSR.
F.5.2. Visualization of Users
In our research, we further conduct a comparative analysis between the user representations in IDSR and MMSR. Specifically, we observed the distribution of these representations within a given user and between different users. The results were illuminating.
The MMSR method reveals a tendency for tighter clustering of diverse items associated with individual users. This characteristic suggests that MMSR, with its multi-modal data, is proficient at discerning and capturing the transition patterns in a user’s activity. Such an attribute is instrumental in understanding and learning the behavior and preferences of users with fewer interactions or ’cold’ items, thus enhancing the overall recommendation quality. Simultaneously, despite achieving compact clusters for items within individual users, MMSR maintains pronounced distances between different users. This aspect illustrates the model’s capacity to uphold the uniqueness of each user’s preferences and interests, preventing the homogenization of user tastes.
This duality of facilitating personalized recommendations while supporting ’cold’ item learning sets MMSR apart. It leverages multi-modal data, enhancing the richness of representation and ensuring a balance between personalization and generalization, an essential aspect of any successful recommendation system. The efficacy of this approach is visually demonstrated in Figure 9, where the clustering and inter-user distances can be clearly observed.
Figure 9. Visualization of users in IDSR and MMSR